Factor the quadratic functions below to determine the zeros of each function. 3. 49x2 = -21x-2 4. 12x2 + 4x - 8 = 0
Answer #1

49x²= = -212-2 49x²+212 +2=0 49x²+14x+74+2=0 7417x+2)+L(72+2l=0 (72+2) (72+1) =0 70+2=0 ar 7x+1=0 - 2 x= -2 or x = 27 Are 4 1